# Hyunjong Ok1,2  Jaeho Lee1 1POSTECH   2HJ AILAB

### v2

@@
Abstract
Singing voices contain much richer information than common voices, such as diverse vocal and acoustic characteristics. However, existing open-source audio-text datasets for singing voices capture only a limited set of attributes and lacks acoustic features, leading to limited utility towards downstream tasks, such as style captioning. To fill this gap, we formally consider the task of singing style captioning and introduce S2Cap, a singing voice dataset with comprehensive descriptions of diverse vocal, acoustic and demographic attributes. Based on this dataset, we develop a simple yet effective baseline algorithm for the singing style captioning. The algorithm utilizes two novel technical components: CRESCENDO for mitigating misalignment between pretrained unimodal models, and demixing supervision to regularize the model to focus on the singing voice. Despite its simplicity, the proposed method outperforms state-of-the-art baselines.
[Uncaptioned image]S2Cap: A Benchmark and a Baseline for Singing Style Captioning
Hyunjong Ok1,2  Jaeho Lee1 1POSTECH   2HJ AILAB hyunjong.ok@gmail.com, jaeho.lee@postech.ac.kr
1 Introduction
Following the recent progress in text-to-speech modeling, the task of speaking style captioning has received a great deal of attention (Yamauchi et al., 2024; Ando et al., 2024; Zhu et al., 2024). Here, the goal is to generate a text prompt that describes para-/non-linguistic characteristics of the speaker from the given audio clip, such as pitch, volume, or gender. The extracted information can contribute greatly to advancing the state-of-the-art of style-conditioned speech synthesis by providing a useful basis for evaluating and labeling the speech data (Wang et al., 2018; Guo et al., 2023; Yang et al., 2024b; Leng et al., 2024).
How much information can speaking style captioning models capture from the singing voices? Singing voices contain rich information regarding the musical characteristics, such as timbre, tempo, or musical genre, providing valuable information for the synthesis and conversion of singing voices (Luo et al., 2020; Zhang et al., 2024; Wang et al., 2024). However, existing speaking style captioning models have been trained to extract only non-musical characteristics from the audio and thus may fall suboptimal for such a purpose. In fact, there is even a lack of appropriate benchmarks to evaluate the performance of such models. Although some datasets provide text attributes or prompts paired with the singing voices (Wang et al., 2022; Huang et al., 2021; Wang et al., 2024), they are limited in the scale of the dataset, the number of attributes, or the diversity of singers (see Table 1).
Figure 1: A comparison of the proposed singing style captioning with related captioning tasks. We did not include speaking style captioning results, as there is no open-sourced model available for the task.
@@ 1 Introduction
Table 1: A comparison of the constructed S2Cap dataset with related singing voice datasets.
@@
NUS-48E Duan et al. (2013) 12 1.9 -
Opencpop Wang et al. (2022) 1 5.3 -
OpenSinger Huang et al. (2021) 66 50.0 -
M4Singer Zhang et al. (2022) 20 29.8 -
Prompt-Singer Wang et al. (2024) 758 306.9 3
@@
Based on the S2Cap dataset, we establish a novel and robust baseline for the singing style captioning. Similar to conventional audio captioning approaches, the proposed framework leverages pretrained audio encoders and text decoders. To align these models, we develop a new strategy, coined CRESCENDO (Connecting Representations for Efficient Sequence-level Captioning with an End-to-end Transformers), which aligns the features of the audio encoder with an auxiliary text encoder. Furthermore, to let the model focus on the vocal track of the audio, our model is trained using a new loss function that uses demixed vocal audio as an additional source of supervision.
Our contributions can be summarized as follows:
  • We propose a new task called singing style captioning, which aims to generate a text prompt capturing the vocal and musical characteristics of the given audio track.
  • We present S2Cap, a comprehensive open-source dataset for singing style captioning with diverse vocal, musical, and demographic attributes.
  • We establish a strong baseline for the singing style captioning, which includes two novel techniques for addressing the representation misalignment and enhancing model performance through vocal demixing supervision.
Figure 2: An illustration of the data generation pipeline of the S2Cap. We start from a base playlist dataset and collect metadata via web scraping. Then, we partition the audio tracks into multiple segments and extract acoustic information from each segment; through these steps, we collect a total of ten attributes (name, age, genre, tag, volume, pitch, gender, timbre, mood, tempo; marked with a subscript) and the singing style description. Finally, we summarize the collected attributes into a single textual prompt using GPT-4o.
2 Related work
Pretrained language and audio models.
Recent advancements in multimodal learning have leveraged pretrained language and audio models to enhance performance across various tasks (Deshmukh et al., 2023; Chu et al., 2024).
Transformer-based architectures have demonstrated remarkable success in NLP, especially when trained on large-scale corpora. While encoder-only models such as BERT (Devlin et al., 2019) excel in representation learning, they are inherently limited in generative tasks, making them unsuitable for captioning. As a result, encoder-decoder architectures (e.g., BART (Lewis et al., 2020)) and autoregressive decoder-based models (e.g., GPT-2 (Radford et al., 2019)) have become the standard choices for text generation tasks, including multimodal captioning (e.g., image captioning, audio captioning).
Pretrained audio models utilize self-supervised learning techniques that have significantly advanced the field of audio representation learning. Models such as Wav2vec 2.0 (Baevski et al., 2020) have been pretrained on large-scale speech datasets to capture rich acoustic features.
For our model and baselines, we construct and integrate audio encoders with text decoders, demonstrating the potential of cross-modal knowledge transfer for improved caption generation.
@@ 2 Related work
The captioning task aims to generate descriptive texts corresponding to an input from diverse modalities.
Various datasets in the visual domain facilitate research for visual captioning. For image captioning, datasets include Flickr30k (Young et al., 2014), MS COCO (Chen et al., 2015), DenseCap (Johnson et al., 2016), and Conceptual Captions (Sharma et al., 2018), each offering diverse annotations for image descriptions. In video captioning, datasets such as MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and LSMDC (Rohrbach et al., 2017) support research of textual descriptions from sequential visual content.
In the auditory domain, datasets such as AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) have been used for general audio captioning tasks. Recently, there has been growing interest in speech style captioning, which involves describing characteristics of human speech (e.g., gender, age group, vocal range) using textual descriptions. PromptTTS (Guo et al., 2023) and InstructTTS (Yang et al., 2024b), for instance, construct dedicated datasets to learn rich speaker-style representations and employ Transformer-based architectures to integrate these representations into synthesized speech. PromptVC (Yao et al., 2024) focuses on text-prompted voice conversion, and PromptSpeaker (Zhang et al., 2023) aims to generate speaker embeddings from text prompts. StyleCap (Yamauchi et al., 2024) initiates speech style captioning to get style descriptions from speech.
@@
Music source separation (MSS), also known as demixing, aims to decompose a mixed music audio signal into its constituent sources, such as vocals, drums, and other accompaniment. Recent advances in deep learning have significantly improved MSS performance, leveraging architectures such as CNN, RNN, and transformer-based networks. Band-split RNN (Luo and Yu, 2023) and HT Demucs (Rouard et al., 2023) have emerged as state-of-the-art models. Band-split RNN employs an RNN-based approach that processes different frequency bands separately, enabling effective modeling of harmonic structures and improving the accuracy of MSS. HT Demucs utilizes a transformer-based architecture to enhance separation performance, capturing local and global dependencies within the audio signal.
In our work, we utilize HT Demucs to extract vocal WAV files from music tracks, which serve as input for our data generation and method.
3 [Uncaptioned image]S2Cap
We now introduce S2Cap, a singing style captioning dataset. The S2Cap consists of 12,105 music tracks with 71,215 textual captions that describe singing styles; each track is partitioned into multiple segments, on which we put separate captions. Along with the captions, each segment has annotations on ten attributes (name, age, gender, tags, volume, pitch, gender, timbre, mood, and tempo). The detailed statistics are given in Table 2.
To avoid license issues, we do not directly share the WAV files for the audio tracks. Instead, we provide the code and URLs to download the wav files. More detailed information and statistics about the dataset are in Appendix A.
Table 2: Basic statistics of the S2Cap dataset
Splits # Tracks # Captions # Words/Caption # Total tokens # Duration (h)
@@
3.1 Construction Pipeline
For the sake of scale, the S2Cap dataset has been constructed by processing an existing musical dataset with an LLM-based pipeline to generate corresponding textual captions (see Fig. 2). As the base dataset, we have used the Melon playlist dataset, a public dataset containing Mel-spectrogram and textual metadata with over 600,000 tracks (Ferraro et al., 2021). Each track in the dataset has been processed as follows.
@@ 3.1 Construction Pipeline
First, we collect textual metadata for each track. This is done by scraping from various web sources, e.g., Last.fm tags. Then, we preprocess the audio track and the metadata as follows. (1) As the Mel-spectrogram in the base dataset is of low resolution, we collect higher-quality audio tracks from YouTube. (2) We filter out the audio files with missing or mismatched metadata entries; we have also filtered out the songs from singers born before 1970, as their songs tend to be missing on YouTube.
@@
Next, we extract acoustic attributes (volume, pitch, gender, timbre, mood, tempo) and style description prompts from each audio segment. This is done with two pretrained models: Qwen-2 Audio (Chu et al., 2024), and a spoken language understanding (SLU) model. [^2] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder Qwen-2 is used to generate annotations on four attributes (gender, timbre, mood, tempo) and the singing style description text; gender and timbre are extracted from the vocal-only version of the audio, and others are generated using the original version. SLU is used to generate volume and pitch annotations. Here, both attributes are classified into three categories—“low,” “medium,” and “high”—where the volume is determined based on the root-mean square of the amplitude and the pitch is determined based on the average $F_{0}$.
Figure 3: An overview of the proposed baseline. An end-to-end transformer comprises an AST encoder and a T5 decoder.
@@
Finally, the extracted attributes and singing style description are summarized into a single textual prompt (per segment), with GPT-4o333in particular, the GPT-4o-2024-08-06 (Hurst et al., 2024). The input prompt we used is illustrated in Appendix C. The generated prompts are post-processed, using num2words [^4] https://github.com/savoirfairelinux/num2words to convert numbers to words and languagetool [^5] https://github.com/languagetool-org/languagetool to correct grammatical errors.
@@
See Appendix B for more details, with an illustrative example of how an audio track is processed with our pipeline.
4 Method
We now describe the proposed baseline method for singing style captioning.
Architecture.
In a nutshell, the model consists of a pretrained audio encoder followed by a pretrained text decoder, which is fine-tuned using the proposed S2Cap dataset (see Fig. 3).
More concretely, consider a Mel-spectrogram of an audio track $X\in\mathbb{R}^{T\times F}$ with $T$ temporal frames and $F$ frequency bins. The audio encoder maps this representation into a feature space
$Z=E_{\text{audio}}(X),$ (1)
where $E_{\text{audio}}(\cdot)$ denotes the audio encoder. In particular, we use the AST encoder (Gong et al., 2021). Given this feature $Z$, the text decoder transforms it into a textual sequence $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{N}]$, which is a sentence that describes the singing style. That is, the text decoder conducts:
$\hat{Y}=D_{\text{text}}(Z),$ (2)
where $D_{\text{text}}(\cdot)$ denotes the text decoder. We use the text decoder that has a paired text encoder $E_{\text{text}}(\cdot)$; in particular, we use T5 (Raffel et al., 2020).
Training.
The model is trained in two steps:
  1. 1.
    Aligning the feature spaces of the audio encoder and text decoder, using the newly proposed method, CRESCENDO (Section 4.1).
  2. 2.
    Jointly fine-tuning the encoder and decoder, with a novel demixing supervision (Section 4.2).
4.1 Aligning with CRESCENDO [Uncaptioned image]
While both pretrained models $E_{\text{audio}},D_{\text{text}}$ are effective when employed separately for each modality, their embeddings may not be aligned well. Indeed, $D_{\text{text}}$ has been trained to process the features generated by a paired $E_{\text{text}}$ (a T5 encoder, in this case), not $E_{\text{audio}}$ (an AST encoder).
To tackle this issue, the proposed CRESCENDO takes a sequence-level similarity learning approach (Grill et al., 2020; Xu et al., 2022). That is, we fine-tune the audio encoder so that their embeddings match the features of the text encoder $E_{\text{text}}$. A technical obstacle here is that the sequence lengths of the audio and textual embeddings are not necessarily identical; we resolve this issue by applying the mean pooling on the audio features (Fig. 4).
More concretely, let $X$ denote the audio drawn from the S2Cap dataset and $Y$ denote the paired textual prompt. We feed each input to the audio and text encoder, respectively, to get the features
$Z_{\text{a}}=E_{\text{audio}}(X),\quad Z_{\text{t}}=E_{\text{text}}(Y).$ (3)
Then, we maximize the cosine similarity between the text features and the mean-pooled audio features. We also apply the stop-grad operation to the $Z_{\text{t}}$, in order to update only the audio encoder. That is, we use the following loss:
$\mathcal{L}_{\text{CRES}}=1-\frac{\text{pool}(Z_{\text{a}})^{\top}\text{sg}(Z_{\text{t}})}{\|\text{pool}(Z_{\text{a}})\|_{2}\cdot\|\text{sg}(Z_{\text{t}})\|_{2}},$ (4)
where $\text{pool}(\cdot),\text{sg}(\cdot),\|\cdot\|_{2}$ denote the mean pooling, stop-gradient, and $\ell_{2}$ norm, respectively.
Figure 4: Align the audio encoder’s latent space with the text encoder through a sequence-level similarity learning technique. A stop-gradient operation is applied to the text encoder.
4.2 Fine-tuning with demixing supervision
After aligning the feature space of the audio encoder with the feature space of the text decoder, we proceed to fine-tuning the whole model on the S2Cap dataset. The training is done with a mixture of two loss functions, which we describe below.
Cross-entropy loss.
First, we use the standard cross-entropy loss for a more accurate generation of textual captions. We adopt the teacher-forcing strategy, i.e., the model receives the true output token from the previous time step as an input for the current time rather than its generated output. Concretely, the loss can be written as:
$\mathcal{L_{\text{CE}}}=-\sum_{n=1}^{N}\log P(y_{n}|y_{1},\ldots,y_{n-1},E_{\text{audio}}(X))$ (5)
where $P(\cdot|\cdot)$ denotes the output probability of the text decoder and $Y=[y_{1},y_{2},\ldots,y_{N}]$ denotes the text prompt paired with audio $X$.
Demixing supervision loss.
The second loss encourages the audio encoder to focus on extracting features from the singing voice itself rather than the sound from the background, such as instrumental accompaniment. For this purpose, we newly developed the demixing supervision, which regularizes the features extracted from the original tracks to be similar to the features from the vocal-only tracks. More concretely, let $X_{\text{voc}}$ be the vocal-only audio from the S2Cap dataset, paired with the original audio $X$. We minimize the Kullback-Leibler (KL) divergence between the features of two audio tracks. That is, we use the loss
$\mathcal{L_{\text{demix}}}=D_{\text{KL}}\left(E_{\text{audio}}(X_{\text{voc}})\|E_{\text{audio}}(X)\right),$ (6)
where $D_{\text{KL}}(\cdot\|\cdot)$ denotes the KL divergence between two embedding distributions.
Final loss. The final loss is the mixture of the cross-entropy and the demixing supervision losses:
$\mathcal{L_{\text{final}}}=\mathcal{L_{\text{CE}}}+\lambda\cdot\mathcal{L_{\text{demix}}},$ (7)
where $\lambda$ is the weight hyperparameter. We have simply used $\lambda=1$ throughout all experiments.
Table 3: Experiment results on the S2Cap test set of our model structure with various audio encoders with T5. For all metrics, the higher, the better.
Encoder Params BLEU4 METEOR ROUGE-$L$ CIDEr SPICE SPIDEr Sentence-BERT
w/ audioset
AST 86M 29.12 31.50 56.25 105.19 43.24 74.21 85.32
w/ music
MERT 95M 21.78 25.30 49.88 58.99 34.16 46.58 81.62
w/ speech
Wav2vec 2.0 base 95M 28.83 31.12 56.37 102.77 44.57 73.67 86.25
HuBERT base 95M 27.64 30.78 55.67 98.07 43.21 70.64 85.52
5 Experiments
@@
We establish comprehensive baselines for our task. We evaluate transformer-based architectures, aligning with our proposed methodology’s framework. We conduct an extensive ablation study across various audio encoder and text decoder combinations. We evaluate five pre-trained models for audio encoding: AST, MERT (LI et al., 2024), Wav2vec 2.0, and HuBERT (Hsu et al., 2021). These are systematically paired with two decoder variants: GPT-2 and BART-base (w/ decoder part only) (Raffel et al., 2020), exploring eight encoder-decoder configurations. Also, we include two specialized models, Prefix-AAC (Kim et al., 2023) and LP-MusicCaps (Doh et al., 2023). They represent the state-of-the-art in related tasks: audio captioning and music captioning.
Table 4: Experiment results on the S2Cap test set of our method and baselines. For all metrics, the higher, the better.
@@
Wav2vec 2.0-GPT2 25.99 30.43 52.41

### v3

# S2Cap: A Benchmark and a Baseline for Singing Style Captioning

@@
Abstract.
[…]
Singing style captioning, Dataset pipeline, Audio-to-text model
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761614submissionid: 1328ccs: Computing methodologies Artificial intelligence
1. Introduction
Following the recent progress in text-to-speech modeling, the task of speaking style captioning has received a great deal of attention (Yamauchi et al., 2024; Ando et al., 2024; Zhu et al., 2024). Here, the goal is to generate a text prompt that describes para-/non-linguistic characteristics of the speaker from the given audio clip, such as pitch, volume, or gender. The extracted information can contribute greatly to advancing the state-of-the-art of style-conditioned speech synthesis by providing a useful basis for evaluating and labeling the speech data (Wang et al., 2018; Guo et al., 2023; Yang et al., 2024a; Leng et al., 2024).
How much information can speaking style captioning models capture from the singing voices? Singing voices contain rich musical characteristics, such as timbre, tempo, or musical genre, providing valuable information for the synthesis and conversion of singing voices (Luo et al., 2020; Zhang et al., 2024). However, existing speaking style captioning models have been trained to extract only non-musical characteristics from the audio and thus may fall suboptimal for such a purpose. In fact, there is even a lack of appropriate benchmarks to evaluate the performance of such models. Although some datasets provide text attributes or prompts paired with the singing voices (Wang et al., 2022; Huang et al., 2021; Wang et al., 2024), they are limited in the scale of the dataset, the number of attributes, or the diversity of singers (see Table 1).
@@ 1. Introduction
Table 1. A comparison of the constructed S2Cap dataset with related singing voice datasets.
@@
NUS-48E (Duan et al., 2013) 12 1.9 -
Opencpop (Wang et al., 2022) 1 5.3 -
OpenSinger (Huang et al., 2021) 66 50.0 -
M4Singer (Zhang et al., 2022) 20 29.8 -
Prompt-Singer (Wang et al., 2024) 758 306.9 3
@@
Building upon the S2Cap dataset, we provide comprehensive baselines by evaluating combinations of diverse audio encoders and text decoders alongside state-of-the-art in related tasks: audio captioning and music captioning. Additionally, we propose a novel training objective designed to enhance the model’s focus on vocal information, utilizing demixed vocal audio as auxiliary supervision. This offers a path for future work to achieve better performance.
Figure 1. An illustration of the data generation pipeline of the S2Cap. We start from a base playlist dataset and collect metadata via web scraping. Then, we partition the audio tracks into multiple segments and extract acoustic information from each segment; through these steps, we collect a total of ten attributes (name, age, genre, tag, volume, pitch, gender, timbre, mood, tempo; marked with a subscript) and the singing style description. Finally, we summarize the collected attributes into a single textual prompt using GPT-4o.
2. Related work
@@
Captioning task aims to generate descriptive texts corresponding to input from diverse modalities.
Various datasets in the visual domain facilitate research for visual captioning. For image captioning, datasets include Flickr30k (Young et al., 2014), MS COCO (Chen et al., 2015), and Conceptual Captions (Sharma et al., 2018), each offering diverse annotations for image descriptions. In video captioning, datasets such as MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and LSMDC (Rohrbach et al., 2017) support research of textual descriptions from sequential visual content.
In the auditory domain, datasets such as AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) have been used for general audio captioning tasks. Recently, there has been growing interest in speech style captioning, which involves describing characteristics of human speech (e.g., gender, age group, vocal range) using textual descriptions. PromptTTS (Guo et al., 2023) and InstructTTS (Yang et al., 2024a), for instance, construct dedicated datasets to learn rich speaker-style representations and employ Transformer-based architectures to integrate these representations into synthesized speech. PromptVC (Yao et al., 2024) focuses on text-prompted voice conversion, and PromptSpeaker (Zhang et al., 2023) aims to generate speaker embeddings from text prompts. StyleCap (Yamauchi et al., 2024) initiates speech style captioning to get style descriptions from speech.
@@
Music source separation (MSS), also known as demixing, aims to decompose a mixed music audio signal into its constituent sources, such as vocals, drums, and other accompaniment. Recent advances in deep learning have significantly improved MSS performance, leveraging architectures such as CNN, RNN, and transformer-based networks. Band-split RNN (Luo and Yu, 2023) and HT Demucs (Rouard et al., 2023) have emerged as state-of-the-art models. Band-split RNN employs an RNN-based approach processing different frequency bands separately, enabling effective harmonic modeling and improving MSS accuracy. HT Demucs utilizes transformer architecture to enhance separation performance, capturing local and global dependencies within audio signals.
In our work, we utilize HT Demucs to extract vocal WAV files, which serve as input for our data generation and method.
3. [Uncaptioned image]S2Cap
We now introduce S2Cap, a singing style captioning dataset. The S2Cap consists of 12,105 music tracks with 71,215 textual captions that describe singing styles; each track is partitioned into multiple segments, on which we put separate captions. Along with the captions, each segment has annotations on ten attributes (name, age, genre, tags, volume, pitch, gender, timbre, mood, and tempo). The detailed statistics are given in Table 2.
To avoid license issues, we do not directly share the WAV files for the audio tracks. Instead, we provide the code and URLs to download the wav files.
Table 2. Basic statistics of the S2Cap dataset.
Splits Tracks Captions Words/Caption Total tokens Duration (h)
@@
3.1. Construction Pipeline
For the sake of scale, the S2Cap dataset has been constructed by processing an existing musical dataset with an LLM-based pipeline to generate corresponding textual captions (see Fig. 1). As the base dataset, we have used the Melon playlist dataset, a public dataset containing Mel-spectrogram and textual metadata with over 600,000 tracks (Ferraro et al., 2021). Although it originates from a Korean music streaming service, “Melon,” it covers a wide range of international music across diverse genres and artists, beyond Korean songs. Each track in the dataset has been processed as follows.
@@ 3.1. Construction Pipeline
First, we collect textual metadata for each track by scraping from various web sources, e.g., Last.fm tags. Then, we preprocess the audio track and the metadata as follows. (1) Since the Mel-spectrogram in the base dataset is of low resolution to avoid licensing issues, we collect higher-quality audio tracks from YouTube. (2) We filter out the audio files with missing or mismatched metadata entries; we also filtered out the songs from singers born before 1970, as their songs tend to be missing on YouTube.
@@
Next, we extract acoustic attributes (volume, pitch, gender, timbre, mood, tempo) and style description prompts from each audio segment. This is done with two pretrained models: Qwen-2 Audio (Chu et al., 2024), and a spoken language understanding (SLU) model. [^2] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder Qwen-2 Audio is used to generate annotations on four attributes (gender, timbre, mood, tempo) and the singing style description text; gender and timbre are extracted from the vocal-only version of the audio, and others are generated using the original version. SLU is used to generate volume and pitch annotations. Here, both attributes are classified into three categories—“low,” “medium,” and “high”—where the volume is determined based on the root-mean square of the amplitude and the pitch is determined based on the average $F_{0}$.
@@
Finally, the extracted attributes and singing style description are summarized into a single textual prompt (per segment), with GPT-4o333in particular, the GPT-4o-2024-08-06 (Hurst et al., 2024).
@@
[…]
4. Experiments
@@
We establish comprehensive baselines for our task. We evaluate transformer-based architectures, aligning with our proposed methodology’s framework. We conduct an extensive ablation study across various audio encoder and text decoder combinations. We evaluate four pretrained models for audio encoding: AST, MERT (LI et al., 2024), Wav2vec 2.0, and HuBERT (Hsu et al., 2021). These are systematically paired with two decoder variants: GPT-2 and BART-base (w/ decoder part only) (Raffel et al., 2020), exploring eight encoder-decoder configurations. Also, we include two specialized models, Prefix-AAC (Kim et al., 2023) and LP-MusicCaps (Doh et al., 2023). They represent the state-of-the-art in related tasks: audio captioning and music captioning. All models are finetuned for $20$ epochs with a batch size of $32$, accumulation steps of $2$, weight decay of $2$, and learning rate of $2\times 10^{-5}$. We use beam search with beam size $5$ during inference.
Evaluation.
To evaluate our proposed methods, we employ metrics that are widely used in captioning tasks such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), SPICE (Anderson et al., 2016), and SPIDEr (Liu et al., 2017). BLEU is a modified n-gram precision metric incorporating a brevity penalty, while ROUGE-L calculates an F-measure based on the longest common subsequence. METEOR enhances the evaluation by considering several factors like stem- and word-level overlap and synonymy. CIDEr employs a geometric mean of n-gram and cosine similarity scores. SPICE focuses on semantic content by parsing scene graphs from captions, and SPIDEr is the average score of SPICE and CIDEr. While the above metrics are valuable for assessing captioning systems, inspired by recent findings (Zhou et al., 2022; Bhosale et al., 2023), they have limits in capturing the semantic meaning in generated captions. So we supplement our evaluation with Sentence-BERT (Reimers and Gurevych, 2019), metrics tailored for improved semantic relatedness, which produces embeddings for calculating sentence-level similarity.
Additional demixing supervision method.
To enhance the model’s ability to effectively represent singing voices against background music, we introduce a novel fine-tuning strategy that incorporates a demixing supervision loss. This approach regularizes the audio encoder to focus on vocal signals. The training is done with a mixture of two loss functions, which we describe below.
Our training objective combines cross-entropy loss for caption generation with our novel demixing supervision loss. The cross-entropy loss employs teacher-forcing strategy, where the model receives the ground-truth token from the previous step as input:
(1) $\mathcal{L_{\text{CE}}}=-\sum_{n=1}^{N}\log P(y_{n}|y_{1},\ldots,y_{n-1},E_{\text{audio}}(X))$
where $P(\cdot|\cdot)$ denotes the output probability of the text decoder and $Y=[y_{1},y_{2},\ldots,y_{N}]$ denotes the text prompt paired with audio $X$. The demixing supervision loss encourages the audio encoder to extract similar features from both original tracks and their vocal-only counterparts by minimizing the KL divergence:
(2) $\mathcal{L_{\text{demix}}}=D_{\text{KL}}\left(E_{\text{audio}}(X_{\text{voc}})\|E_{\text{audio}}(X)\right),$
where $D_{\text{KL}}(\cdot\|\cdot)$ denotes the KL divergence between two embedding.
The final loss is the mixture of the cross-entropy and the demixing supervision:
(3) $\mathcal{L_{\text{final}}}=\mathcal{L_{\text{CE}}}+\lambda\cdot\mathcal{L_{\text{demix}}},$
where $\lambda$ is the weight hyperparameter.
Table 3. Experiment results on the S2Cap test set of our method and baselines. For all metrics, the higher, the better.
@@
Wav2vec 2.0-GPT2 25.99 30.43 52.41 94.37 44.22 69.30 85.01
HuBERT-GPT2 25.80 30.56 52.45 95.41 44.62 70.01 85.28
w/ BART
AST-BART 26.47 30.38 52.50 99.42 43.98 71.70 84.31
MERT-BART 26.49 30.43 52.50 99.95 43.92 71.93 84.04
Wav2vec 2.0-BART 25.51 30.01 51.97 93.30 43.81 68.56 84.53
HuBERT-BART 26.18 30.27 52.60 96.45 44.34 70.39 84.76
w/ T5
AST-T5 29.12 31.50 56.25 105.19 43.24 74.21 85.32
MERT-T5 21.78 25.30 49.88 58.99 34.16 46.58 81.62
Wav2vec 2.0-T5 28.83 31.12 56.37 102.77 44.57 73.67 86.25
HuBERT-T5 27.64 30.78 55.67 98.07 43.21 70.64 85.52
SOTA of related works
Prefix-AAC 29.47 31.64 56.84 104.10 44.87 74.48 86.38
+ w/ Demixing supervision 29.70 31.74 56.97 105.70 44.89 75.29 86.45
LP-MusicCaps 28.33 31.06 55.60 102.92 42.61 72.77 84.91
Figure 2. An example from S2Cap. The emotional intensity of the music gradually escalates from the first to the latter half, effectively captured in the generated caption.
5. Results
5.1. Experiment results
Experiments in various baselines.
We conducted experiments using various encoder-decoder combinations and state-of-the-art (SOTA) models from related works, which are shown in Table 3. Additionally, we applied demixing supervision to the best-performing model Prefix-AAC from our baseline experiments, demonstrating improved performance with demixing supervision.
5.2. Data quality assessment
Human evaluation.
To assess the quality of our dataset, we conducted a human evaluation study by comparing captions generated by GPT-4o, Qwen2.5-72B-Instruct (Yang et al., 2024c), and Llama3.3-70B-Instruct (Dubey et al., 2024) models. Three human annotators were given 200 sampled captions with audio and asked to determine which model produced the best outputs. As shown in Fig. 3, GPT-4o consistently outperformed other models, demonstrating superior caption quality.
In addition to comparative evaluation, we further analyzed GPT-4o-generated caption quality by assessing consistency and fluency criteria, scored out of 5 and 3, respectively. Consistency was evaluated based on factual alignment with music audio tracks, while fluency was measured in terms of grammar, spelling, punctuation, word choice, and sentence structure. Three human annotators were given the same 200 sampled captions with audio and asked to evaluate caption quality. As shown in Table 4, GPT-4o achieved an average consistency score of 4.94 and fluency score of 2.97, confirming the high quality of its generated captions.
To check the objectivity, we have conducted a human evaluation of the timbre generated by Qwen-2 Audio, where the subjective terms come from. Specifically, 20 annotators judged whether the generated timbre appropriately matched one of the 20 well-known singers. Timbre accuracy is 0.75, indicating a strong alignment between human perception and Qwen-2 Audio. Recent studies report human-LLM alignment in evaluations typically about 70% (Kolchinski et al., 2019; Yang et al., 2024b), our result demonstrates notably high agreement.
Table 4. The result of human evaluation on the quality of captions generated by S2Cap.
Methods Consistency Fluency Timbre Acc.
Human eval 4.94 2.97 0.75
Figure 3. Human evaluation results comparing generated captions. The numbers in the bar plot indicate the win rate of the model on the left.
Captioning examples.
We show an example of our S2Cap dataset, as illustrated in Fig. 2. These captions effectively capture various attributes of singing, demonstrating high-quality caption generation. In the given example song, the verse part features a soft and emotional vocal, whereas the chorus gradually intensifies, culminating in an explosive emotional peak. Our dataset successfully reflects these dynamic shifts in the same song.
6. Conclusion
We propose a novel task, singing style captioning, which aims to generate textual prompts describing the vocal characteristics of singers from given song inputs. For this task, we developed S2Cap, a comprehensive dataset reflecting diverse vocal attributes, and established a robust baseline method that effectively captures singing voice characteristics. These contributions provide a solid foundation for future research in this emerging field.
Usage of Generative AI
We use for manuscript refinement and code optimization. Additionally, LLMs are utilized into our data generation pipeline.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00453301), and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).